Solution A is 25% acid and solution B is 40% acid. How much of each solution is needed to make 100 liters of 30% acid?

Answer by stanbon(75887)

Solution A is 25% acid and solution B is 40% acid. How much of each solution is needed to make 100 liters of 30% acid?
-------------
One variable solution:
0.25x + 0.40(100-x) = 0.30*100
Multiply thru by 100 to get:
25x + 40*100 - 40x = 30*100
-15x = -10*100
x = (2/3)100
x = 66.67 liters (amt of 25% solution needed in the mixture)
---
100-x = 33.3 liters (amt of 40% solution needed in the mixture)
==================================
Cheers,
Stan H.